how do I solve x + 7y = 39 and 3x - 2y = 2 by using substitution?

Answer by checkley79(3341)

x + 7y = 39
X=39-7Y
3x - 2y = 2
3(39-7Y)-2Y=2
117-21Y-2Y=2
-23Y=2-117
-23Y=-115
Y=-115/-23
Y=5 ANS.
X+7*5=39
X+35=39
X=39-35
X=4 ANS.
PROOF:
3*4-2*5=2
12-10=2
2=2